Suppose the average consumer divides up his expenditure between bread, meat, milk and cigarettes in the ratio 4:3:2:1. During the course of a year, the price of bread falls by 10% while that of meat rises by 20% and those of milk and cigarettes both increase by 10%. What is the increase in the price level during the year?
1
Expert's answer
2020-04-28T07:54:20-0400

Base expenditures:

"4U+3U+2U+1U=10"  Units

New expenditures:

"4*0.9U+3*1.2U+2*1.1U+1*1.1U=10.5U"

Change in price levels:

"(10.5U-10U)\/10U=5" %